A production manager is tasked to determine which one of 2 suppliers from whom to source. However, each supplier may not be able to reliably fulfill the firm’s orders on a consistent basis. The manager estimates, through some research, that if supplier 1 is chosen, the probability that they will be unable to fulfill orders is 0.3, but it would cost the manager’s firm $1,000,000 in this instance. On the other hand, if supplier 1 is able to fulfill the orders with a probability of 0.7, the firm will be able to profit $500,000. Likewise, if supplier 2 is chosen, the probability that they will be unable to fulfill orders is 0.6, but it would cost the manager’s firm $300,000 in this instance. On the other hand, if supplier 2 is able to fulfill the orders with a probability of 0.4, the firm will be able to profit $600,000. If the manager does not make a decision, they will earn a profit of $0.

What is the logical solution for the manager to choose? (a) Contract with supplier 1 (b) Contract with supplier 2 (c) Do not contract at all (d) Split the order between supplier 1 and supplier 2 evenly (that is, have 50% of orders go to supplier 1, and 50% of orders go to supplier 2)
Answer #1

If Supplier 1 is chosen, EMV(Supplier 1) = 0.3*(-1000000)+0.7*500000 = 50000

If Supplier 2 is chosen, EMV(Supplier 2 ) = 0.6*(-300000)+0.4*600000 = 60000


EMV(No decision) = 0

hence the logical solution that the manager should chose b) Contract with Supplier 2 as this solution has highest EMV